OpenStudy (anonymous):

George bought a shirt and a pair of pants at a clothing store. The pants cost $5.91 more than the shirt, and he used a $15 off coupon to purchase the shirt. The equation below is used to determine the cost of the shirt before the coupon, s. s - $15 = $15.82 How much did the pants cost? $36.73 $21.73 $24.91 $30.82

OpenStudy (the_fizicx99):

D gurrrrl :> S - 15 = 15.82 +15 +15 = 30.82